Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Monday, November 9, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 9, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/09/98 (Monday)

Segment 1 (Emidio to ANF): Report for 11/09/98. The Hydrostatic test for the 8 - Mile pipe segment that extends from EMP 33.34 to EMP 25.46 took place today. An excavator potholed the ROW on the south side of the large culverts at EMP 25.55 to test the soil and backfilled the hole at the end of the day. At Lebec road (~EMP 25.20), trench was excavated for a tie-in to the directional drill pipe section. A small section of trench was opened on the Tejon Bench, where the previously buried pipe segment's end was exposed for tie-in work at EMP 18.49. Immediately north, the pipe was padded and backfilled slowly, by loader because of the narrow work area. An excavator continued to open trench along Grapevine road, working north in the area of EMP 15.98. Station work continued at both Grapevine and Emidio. A long pipe segment was lowered in to trench along Aqueduct Access Road B, near EMP 6.06. Padding of the pipe took place along the same road to the east near EMP 6.50. Cutout and repair work was done on the welded pipe segment along Hwy 166 at ~EMP 0.3 while shoring /dewatering of the trench continued to the east along the north shoulder of the highway. Welding and coating of pipe continued at the bore site along Rock road at EMP 8.96.

Segment 2 (ANF to Sylmar) Report for 11/09/98: 2A (Templin Hwy to Castaic): Work began on the tie-in of the testing cutout at Templin Hwy, EMP 50.89 to 50.91. Construction of the block wall on the ANF access road south of Templin continued. Revegetation continued between stream crossing #46 and the I-5 bore. Erosion prevention work occurred south of the I-5 bore at EMP 53.59. Coating and trenchbreaker installation progressed north to EMP 55.64. Shading reached 55.72. F/O conduit laying and backfilling reached EMP 55.84. Welders worked on the anchor structure at the Lake Hughes block valve, EMP 56.62. 2B (Castaic to Sylmar): F/O cable was pulled between Lake Hughes Rd and Church Street. Boring of the Feedmill Rd directional drill reached 40 feet, then struck a gravel layer. This section may be redone. A frac tank was hauled in for disposal of excess drilling mud. At Calgrove, backfill of both bore pits continued. Cutting of concrete roadway north of Sierra Hwy continued. South of Sierra Hwy, 200 feet of trench was excavated. Between Sierra Hwy and Sunshine Summit, pipe-lowering, welding, and backfilling occurred. 150 feet of asphalt was applied under the I-5 bridge at Sunshine Summit.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): Report for 9 November 1998: 3A: Backfill crews worked at final clean up and grading at Verdugo Wash today. That crew also pulled fiber optic conduit through the casing on the Verdugo Wash Bridge. Electrical work occurred at the block valve just south of Buena Vista Avenue in Burbank. The crew at the Tuxford Street bore site continued to secure the bore pit. 3B: No work occurred. 3C: The bore crew at Hubbard spent the day pulling, then replacing, their augers in order to verify proper alignment. Backfill continued to work on the ROW between Roxford Street and Grapevine Creek. A small crew worked on clean up of the forms at the California Aqueduct crossing. Welding and x-ray worked late on the pipe for the aqueduct crossing.

Segment 4B (El Segundo to Lynwood) 11-9-98: Backfilling activities continued at Santa Ana and Alameda ESEMP 0.04. Construction activities continued at El Segundo station ESEMP 11.00. This crew was pushing the pipe through the casing. The crew at Slater ESEMP 1.68 was testing the fiberoptic conduit. The crew at Lynwood station EMP 0.0. was dewatering today. backfilling activities continued at Western bore pit ESEMP 5.56.

Segment 5 (Lynwood to Long Beach) Report for 9 November 1998: The horizontal directional drill at Highway 91/Compton Creek experienced a frac-out at about 2:30 p.m. The mud was contained and work was shutdown until a vacuum truck could arrive. At the Wilmington refinery, the crew continues to push pipe through the casing without problems.

Inspection Summary

In the past 24 hours, 5 inspection event(s) were reported for 5 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page